OpenStudy (mony01):

A=4i+3j B=2i+5j What is the equation of the resultant vector? What angle does it make with respect to the x axis?

OpenStudy (fibonaccichick666):

so, resultant means what in this case? are we adding, subtracting, multiplying?

OpenStudy (mony01):

C?

OpenStudy (fibonaccichick666):

...? that makes no sense

OpenStudy (mony01):

I think when we graph those vectors there's a resultant which would be C

OpenStudy (mony01):

so adding?

OpenStudy (zzr0ck3r):

you are taking two object and talking about a result. Well, we need to know for what operation. I am guessing that we are talking about vector addition, but maybe not?

OpenStudy (zzr0ck3r):

If, indeed they mean addition, then you add the like parts. \(A+B = 4i+3j+2i+5j=4i+2i+3j+5j=(4+2)i+(3+5)j=6i+8j\). To get the angle we note that \(\tan(\theta)=\frac{\text{rise}}{\text{run}}=\frac{8}{6}\implies \theta = \tan^{-1}(\frac{4}{3})\approx53.13^{\circ}\)